# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ C.1 Approximation Error
$\mathbb{E}\|\mathbf{x}_{t}\|^{4}$ $\leq 8\mathbb{E}\|\mathbf{x}_{1}\|^{4}+8\sigma^{4}\mathbb{E}\|\mathbf{B}_{1-t}\|^{4}$
$\leq 8{d^{*}}^{2}+8\sigma^{4}(1-t)^{2}d^{*}(d^{*}+2).$
Therefore,
$\mathbb{E}\|\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{4}\leq\frac{8(9{d^{*}}^{2}+8\sigma^{4}(1-t)^{2}d^{*}(d^{*}+2))}{\sigma^{8}(1-t)^{4}}.$
Then, we estimate $\mathbb{P}(\|\mathbf{x}_{t}\|_{\infty}>r)$. It holds that
$\mathbb{P}(\|\mathbf{x}_{t}\|_{\infty}>r)$ $=\mathbb{P}(\|\mathbf{x}_{1}+\sigma\mathbf{B}_{1-t}\|_{\infty}>r)$
$\leq\sum_{i=1}^{d^{*}}\mathbb{P}(|x_{1,i}+\sigma B_{1-t,i}|>r)$
$\leq\sum_{i=1}^{d^{*}}\mathbb{P}\left(\left|\frac{B_{1-t,i}}{\sqrt{1-t}}\right|>\frac{r-1}{\sigma\sqrt{(1-t)}}\right)$
$\leq 2d^{*}\exp\left({-\frac{(r-1)^{2}}{2\sigma^{2}(1-t)}}\right).$
Thus, the second term can be bounded by
$2\sqrt{2d^{*}}\left(\widehat{K}^{2}+\frac{\sqrt{8(9{d^{*}}^{2}+8\sigma^{4}(1-t)^{2}d^{*}(d^{*}+2))}}{\sigma^{4}(1-t)^{2}}\right)\exp\left({-\frac{(r-1)^{2}}{4\sigma^{2}(1-t)}}\right).$ (13)
By substituting $\widehat{K}$ into (13) and setting it to be less than $\epsilon^{2}$, we obtain
$r=\mathcal{O}\left(\sqrt{\log{\frac{d^{*}}{\epsilon(1-T)}}}\right),$
@@
$\mathbb{E}\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\leq(1+d^{*})\epsilon^{2}.$
Substituting $r$ into network configuration yields that
@@
$L=\mathcal{O}\left(\log{\frac{1}{\epsilon}}+d^{*}\right),M=\mathcal{O}\left(\frac{{d^{*}}^{\frac{5}{2}}\left(\log{\frac{d^{*}}{\epsilon(1-T)}}\right)^{\frac{d^{*}+3}{2}}\xi^{d^{*}}\epsilon^{-(d^{*}+1)}}{(1-T)^{3}}\right),$
@@
$J=\mathcal{O}\left(\frac{{d^{*}}^{\frac{5}{2}}\left(\log{\frac{d^{*}}{\epsilon(1-T)}}\right)^{\frac{d^{*}+3}{2}}\xi^{d^{*}}\epsilon^{-(d^{*}+1)}}{(1-T)^{3}}\left(\log{\frac{1}{\epsilon}}+d^{*}\right)\right),K=\mathcal{O}\left(\frac{\sqrt{d^{*}\log{\frac{d^{*}}{\epsilon(1-T)}}}}{1-T}\right),$

### v2

@@ B.4 Error bound of approximating $\nabla\log q_{t}(\mathbf{x})$ with ReLU neural network
[…]
to approximate $\frac{[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}}{\sigma_{t}},1\leq i\leq d^{*}$.
Now, we analyse the error bound between $[\mathbf{s}_{\mathrm{score}}^{(1)}]_{i}$ and $\frac{[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}}{\sigma_{t}},1\leq i\leq d^{*}$. The approximation error can be bounded by
$\epsilon_{\mathbf{s}_{\mathrm{score}}^{(1)},i}\leq\epsilon_{\mathrm{prod},2}+2C_{5}\max\left\{\epsilon_{\mathbf{s}_{6},i},\left|\mathrm{s}_{\mathrm{rec},2}(\mathrm{s}_{\mathrm{root}}(\sigma_{t}))-\frac{1}{\sigma_{t}}\right|\right\},$
where $C_{5}=\max\{\sup_{t}\sigma_{t},\sup_{t,\mathbf{x}}[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}\}=\mathcal{O}(\sqrt{\log\mathcal{R}})$ and $\epsilon_{\mathbf{s}_{6},i}=\sup_{t,\mathbf{x}}\left|[\mathbf{s}_{6}(t,\mathbf{x})]_{i}-[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}\right|$. $\epsilon_{\mathbf{s}_{6},i}$ can be bounded as follows:
$\epsilon_{\mathbf{s}_{6},i}\lesssim\epsilon_{\mathrm{prod},1}+2C_{6}\max\left\{\epsilon_{0}+\epsilon\log^{\frac{d^{*}+1}{2}}\epsilon^{-1},\epsilon_{\mathrm{rec},1}+\frac{\epsilon_{0}+\epsilon\log^{\frac{d^{*}}{2}}\epsilon^{-1}}{\epsilon_{\mathrm{rec},1}^{2}}\right\},$
where $C_{6}\lesssim\max\left\{\sup_{t,\mathbf{x}}|[\mathbf{h}_{1}(t,\mathbf{x})]_{i}|,\sup_{t,\mathbf{x}}|g_{1}(t,\mathbf{x})\vee\mathcal{R}^{-(2\beta+1)}|\right\}=\mathcal{O}(1)$. Similarly, we have
$\left|\mathrm{s}_{\mathrm{rec},2}(\mathrm{s}_{\mathrm{root}}(\sigma^{2}t))-\frac{1}{\sigma_{t}}\right|\leq\epsilon_{\mathrm{rec},2}+\frac{\sigma\epsilon_{\mathrm{root}}}{\epsilon_{\mathrm{rec},2}^{2}}.$
By taking
$\epsilon=\mathcal{R}^{-(3\beta+2)},\leavevmode\nobreak\ \epsilon_{\mathrm{prod},1}=\frac{\mathcal{R}^{-\beta}}{8C_{5}},\leavevmode\nobreak\ \epsilon_{\mathrm{prod},2}=\frac{\mathcal{R}^{-\beta}}{2},\leavevmode\nobreak\ \epsilon_{\mathrm{rec},1}=\frac{\mathcal{R}^{-\beta}}{32C_{5}C_{6}},\leavevmode\nobreak\ \epsilon_{\mathrm{rec},2}=\frac{\mathcal{R}^{-\beta}}{8C_{5}},$
@@
$\epsilon_{0}=\min\left\{\frac{\mathcal{R}^{-\beta}}{16C_{5}C_{6}},\left(\frac{\mathcal{R}^{-\beta}}{32C_{5}C_{6}}\right)^{3}\right\},\leavevmode\nobreak\ \epsilon_{\mathrm{root}}=\frac{1}{\sigma}\cdot\left(\frac{\mathcal{R}^{-\beta}}{8C_{5}}\right)^{3},$
we obtain
$\epsilon_{\mathbf{s}_{6},i}\lesssim\frac{\mathcal{R}^{-\beta}}{C_{5}},\leavevmode\nobreak\ \epsilon_{\mathbf{s}_{\mathrm{score}}^{(1)},i}\lesssim\mathcal{R}^{-\beta}.$
Subsequently, we can obtain the network parameters of $[\mathbf{s}_{\mathrm{score}}^{(1)}]_{i}$, $1\leq i\leq d^{*}$:
$L_{i}=\mathcal{O}(\log^{4}\mathcal{R}),M_{i}=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{7}\mathcal{R}),J_{i}=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{9}\mathcal{R}),\kappa_{i}=\exp\left(\mathcal{O}(\log^{4}\mathcal{R})\right).$
Combining $[\mathbf{s}_{\mathrm{score}}^{(1)}]_{i}$, $1\leq i\leq d^{*}$, and by Lemma F.7, we construct the ReLU neural network
$\mathbf{s}_{\mathrm{score}}^{(1)}:=\left[[\mathbf{s}_{\mathrm{score}}^{(1)}]_{1},[\mathbf{s}_{\mathrm{score}}^{(1)}]_{2},\cdots,[\mathbf{s}_{\mathrm{score}}^{(1)}]_{d^{*}}\right]^{\top},$
with network parameters
$L=\mathcal{O}(\log^{4}\mathcal{R}),M=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{7}\mathcal{R}),J=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{9}\mathcal{R}),\kappa=\exp\left(\mathcal{O}(\log^{4}\mathcal{R})\right),$
and it satisfies $\|\mathbf{s}_{\mathrm{score}}^{(1)}(t,\cdot)\|_{\infty}\lesssim\frac{\sqrt{\log\mathcal{R}}}{\sigma_{t}}$. Thus, the approximation error $(A)$ is bounded by
$(A)\lesssim d^{*}\mathcal{R}^{-2\beta}\lesssim\mathcal{R}^{-2\beta}.$ (16)
Next, we bound the approximation error $(B)$. Recall that $\|\mathbf{x}\|_{\infty}\leq 1+C\sigma_{t}\sqrt{(2\beta+1)\log\mathcal{R}}\leq 1+C\sigma\sqrt{(2\beta+1)\log\mathcal{R}}$ and $q_{t}(\mathbf{x})\geq\mathcal{R}^{-(2\beta+1)}$ hold. In this case, we have $\|\nabla\log q_{t}(\mathbf{x})\|\lesssim\frac{\sqrt{\log\mathcal{R}}}{\sigma_{t}}.$ We first consider the case $\mathbf{x}\in[-1,1]^{d^{*}}$. For $1\leq i\leq d^{*}$, we have
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left|\frac{[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}}{\sigma_{t}}-[\nabla\log q_{t}(\mathbf{x})]_{i}\right|$
$=\left|\frac{[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}}{\sigma_{t}}\right|\left|1-\frac{[\nabla\log q_{t}(\mathbf{x})]_{i}}{[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}/\sigma_{t}}\right|$
$\lesssim\frac{\sqrt{\log\mathcal{R}}}{\sigma_{t}}\frac{|[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}-\sigma_{t}[\nabla\log q_{t}(\mathbf{x})]_{i}|}{|[\mathbf{h}^{\prime}(t,\mathbf{x})]_{i}|}$
@@
$\lesssim\frac{\sqrt{\log\mathcal{R}}}{\sigma_{t}}\frac{\left|\frac{[\mathbf{h}_{1}(t,\mathbf{x})]_{i}}{g_{1}(t,\mathbf{x})\vee\mathcal{R}^{-(2\beta+1)}}-\frac{\sigma_{t}[\nabla q_{t}(\mathbf{x})]_{i}}{q_{t}(\mathbf{x})}\right|}{\left|\frac{[\mathbf{h}_{1}(t,\mathbf{x})]_{i}}{g_{1}(t,\mathbf{x})\vee\mathcal{R}^{-(2\beta+1)}}\right|}$